# Polariton lattices as binarized neuromorphic networks

### v1

@@ Results
In the proposed approach, the preparation of the input layer involves partial overlapping of the initial image with itself. To describe such overlapping, we introduce the parameter of densing degree $s$, which can take both integer and fractional values. For instance, a densing degree of 2 means that each element in the input layer is on average determined by two elements of the initial binarized layer. Similarly, with ${s=1.5}$, approximately half of the input layer elements are on average determined by one element of the initial layer, while the other half by two elements. Figure 5 illustrates how the classification accuracy of the polaritonic network, with different numbers of neurons in the hidden layer, changes with the densing degree $s$ of the input layer. The key finding of our study is that with approximately 1350 neurons (corresponding to ${n_{\text{in}}=51}$) and a densing degree of about 3.25, the accuracy reaches the previously reported benchmark of 96% [12, 13] (see red star on a violet curve in Fig. 5), and surpasses it with the further increase of the number of neurons. Our predicted maximum average accuracy is 97.5%, obtained for the network with the input matrix size of ${n_{\text{in}}=160}$, corresponding to the number of neurons less than $1.3\times 10^{4}$, and the densing degree ${s=3.7}$, see the red star at the brown curve in Fig. 5. The accuracy can be further increased by at least 0.1% through the appropriate choice of a randomization mask. It is evident that the dependencies exhibit a pronounced maximum, and the value of $s$, at which this maximum occurs, increases with the growing number of neurons $N_{\text{d}}$. Moreover, the fewer the number of neurons, the steeper the dependencies become. It can also be observed that the impact of the randomization mask notably decreases with an increase of the number of neurons.
@@ Discussion
Transitioning to the practical aspects of this architecture, it becomes essential to consider the possible experimental realization of polariton condensate lattices designed above. Spatially-distributed systems comprising chains and lattices of interconnected quantum entities have gained recognition as platforms for information storage, transmission and processing, as well as simulators of complex phenomena [17, 37, 38, 39, 19]. Polariton lattices, a recent breakthrough in spatially-distributed quantum systems, are notable for their exceptional spatial coherence [27, 21, 26, 22]. This coherence significantly exceeds that of individual condensates and it facilitates phase locking of nodes across the entire lattice. This widespread phase synchronization could potentially enhance the nonlinearity of interactions essential for neuromorphic network operation, see, e. g., [3]. However, this scenario leads to the loss of a key advantage emphasized in our work: the feasibility of neuron response binarization would be significantly compromised, impacting computational speed and resource efficiency. To address this issue while retaining the advantage of polariton lattices, which is the optical controllability of connections between condensates, we propose ensuring pairwise interactions of condensates within the lattice by selectively severing non-contributing links. For this purpose, a variety of approaches exists.
@@
The network’s effectiveness was demonstrated using the MNIST dataset for handwritten digit recognition. Our network has not only shown competitive performance against existing systems, but also surpassed them when taking advantage of the original signal densing technique. The developed approach allowed the network to achieve a remarkable 97.5% classification accuracy, theoretically.
@@ Methods
$U_{\text{d}}(t,\mathbf{r})=\left(\frac{g_{1}+iR_{1}}{2(\gamma_{\text{R}}+R_{1}|\Psi(t,\mathbf{r})|^{2})}+G_{1}\right)P_{1}(\mathbf{r})$ (3)
@@
characterizes the complex effective potential arising from the excitation of polariton condensates in a dyad through non-resonant optical pumping with intensity $P_{1}(\mathbf{r})$. $g_{1}$ is the constant of interaction of polaritons with reservoir excitons, $R_{1}$ is the stimulated scattering rate from the reservoir to the condensate, $\gamma_{\text{R}}$ is the exciton decay rate. The parameter $G_{1}$ characterizes repulsion from the dark exciton reservoir, which also inevitably emerges within the pump spot.
@@
The third term in Eq. (2) is responsible for the potential, arising from the signal beam of intensity $P_{2}(\mathbf{r})$:
$U_{\text{in}}(t,\mathbf{r})=\left(\frac{g_{2}+iR_{2}}{2(\gamma_{\text{R}}+R_{2}|\Psi(t,\mathbf{r})|^{2})}+G_{2}\right)P_{2}(\mathbf{r}).$ (4)
@@
In Ref. [27], it is suggested to use cross-circularly polarized pump $P_{2}(\mathbf{r})$ in relation to $P_{1}(\mathbf{r})$. This approach aims to reduce gain of the condensates in the dyad due to their overlapping with the signal. The parameters in Eq. (4) for $U_{\text{in}}(t,\mathbf{r})$ have the same meaning as those in Eq. (3) for $U_{\text{d}}(t,\mathbf{r})$. We take the pump intensities in a Gaussian form as $P_{1}(\mathbf{r})\propto\sum_{j,k=\pm}\exp\left\{-[(x+jd/2)^{2}+(y+kd/2)^{2}]/2w_{1}^{2}\right\}$ and $P_{2}(\mathbf{r})\propto\exp\left\{-r^{2}/2w_{2}^{2}\right\}$, where $w_{1,2}$ are the widths of the pumps, $d$ is the distance between the centers of the pump spots.
@@
The last term in Eq. (2), which differs our model from one in Ref. [27] characterizes the stationary barrier. We take it as $V(\mathbf{r})=V_{0}\sum_{j=\pm}\exp\left[-(x+ju/2)^{8}/u_{x}^{8}+y^{4}/u_{y}^{4}\right]$, where $u$ is the distance between the spatial barrier components. The magnitude $V_{0}$ can be both real and imaginary, depending on the nature of the barrier.
@@
For simulation, we take the following parameters suggested in Ref. [27]. The effective polariton mass is ${m^{*}=0.49\,\text{meV ps}^{2}\mu\text{m}^{-2}}$, the decay rates are ${\gamma=1/6\,\text{ps}^{-1}}$ and ${\gamma_{\text{R}}=0.05\,\text{ps}^{-1}}$, the interaction constants are ${\alpha=2.4\,\mu\text{eV}\,\mu\text{m}^{2}}$, ${g_{1}=0.8\alpha}$ and ${g_{2}=1.8\alpha}$, repulsion constants from dark reservoirs are $G_{1,2}=4g_{1,2}/\gamma_{\text{R}}$, the widths of the pump spots are ${w_{1}=1.06\,\mu\text{m}}$ and ${w_{2}=1.6w_{1}}$. We also take the scattering rates as ${R_{1}=7\alpha}$ and ${R_{2}=0.7R_{1}}$, and the distance between the pump spots as ${d=8\,\mu\text{m}}$. For the barrier potential, we take the distance ${u=11\,\mu\text{m}}$ and the widths ${u_{x}=2\,\mu\text{m}}$, ${u_{y}=0.7\,\mu\text{m}}$. Remarkably, the used for simulation pure real (Figs. 1(h,j)) and pure imaginary (Figs. 1(k,l)) barriers possess similar shape and differ in magnitude, which is a control parameter of the model.

### v2

@@ Results
In the proposed approach, the preparation of the input layer involves partial overlapping of the initial image with itself. To describe such overlapping, we introduce the parameter of densing degree $s$, which can take both integer and fractional values. For instance, a densing degree of 2 means that each element in the input layer is on average determined by two elements of the initial binarized layer. Similarly, with ${s=1.5}$, approximately half of the input layer elements are on average determined by one element of the initial layer, while the other half by two elements. Figure 4(c) illustrates how the classification accuracy of the polaritonic network, with different numbers of neurons in the hidden layer, changes with the densing degree $s$ of the input layer. The figure reveals pronounced peaks in the accuracy dependencies. This phenomenon can be explained as follows. States both in the input space and in the feature space are represented by binary vectors consisting of various combinations of ‘0’s and ‘1’s. Increasing the densing degree $s$ implies a higher fraction of ‘1’s in the input space, which in turn activates more neurons in the hidden layer. This activation facilitates structural nonlinearity through the network’s architecture, thereby enhancing the distinct features of elements belonging to specific data classes (digits in our case). Consequently, this also increases the number of ‘1’s in the feature space.
However, beyond a certain point, further increasing the densing degree $s$ results in diminishing returns. Additional ‘1’ in the feature space no longer contribute to highlighting distinctive features. Instead, they may obscure these features by overwhelming other state elements or adding redundant information. In the extreme case where $s$ approaches infinity, every input state would be mapped to a feature space vector consisting entirely of ‘1’s, rendering state recognition impossible. For a given number of involved neurons, the peak accuracy is achieved when the number of ‘0’s and ‘1’s is comparable with each other. The variability in the peak’s position and the rate at which classification accuracy changes with different numbers of neurons in the hidden layer can be attributed to the structural nonlinearity introduced by the random pairing of input signals with neurons. This nonlinearity varies with each configuration, significantly impacting how effectively the network processes and utilizes the increased density of ‘1’s to distinguish between classes. Consequently, the network’s ability to manage this increased input density depends on the specific configuration and number of neurons.
The key finding of our study is that with approximately 800 neurons (corresponding to ${n_{\text{in}}=39}$) and a densing degree of about 3, the accuracy reaches the previously reported benchmark of 96% [15, 16] (see red marker on a red curve in Fig. 4(c)), and surpasses it with the further increase of the number of neurons. Our predicted maximum average accuracy is 97.5%, obtained for the network with the input matrix size of ${n_{\text{in}}=160}$, corresponding to the number of neurons less than $1.3\times 10^{4}$, and the densing degree ${s=3.5}$, see the red marker at the brown curve in Fig. 4(c). The accuracy can be further increased by at least 0.1% through the appropriate choice of a randomization mask. It is evident that the dependencies exhibit a pronounced maximum, and the value of $s$, at which this maximum occurs, increases with the growing number of neurons $N_{\text{d}}$. Moreover, the fewer the number of neurons, the steeper the dependencies become. It can also be observed that the impact of the randomization mask notably decreases with an increase of the number of neurons.
Figure 5: Evaluation of the Speech Commands [37, 38] recognition (ten commands) by the polariton neuromorphic network. (a) The recognition accuracy in dependence on the size of the input signal lattice $n_{\text{in}}$ (lower scale) or the number of neurons (dyads) in the hidden layer $N_{\text{d}}$ (upper scale). (c) The recognition accuracy in dependence of the overlap degree $s$ of the input signal for square polariton lattice systems of different size $d_{\text{in}}$ with different numbers of neurons in the interaction layer: $d_{\text{in}}=20$ with 220 neurons (blue), $d_{\text{in}}=28$ with 420 neurons (green), $d_{\text{in}}=39$ with 800 neurons (red), and $d_{\text{in}}=160$ with 12960 neurons (brown). Red markers indicate the maxima of the dependencies. Horizontal dashed lines indicate the accuracy levels for alternative classification approaches (from bottom to top): the linear software classification of the binarized (49.2%) and non-binarized (61%) MFCC feature matrices, and the HMM-GMM-based classification of the MFCC feature matrices (65.8%).
Speech Commands dataset analysis
To further evaluate the adaptability of our proposed neuromorphic network and its ability to generalize across different types of data beyond the MNIST dataset, we tested it using the Speech Commands dataset [37, 38]. This dataset fundamentally differs from MNIST in structure, consisting of one-second audio clips containing spoken words rather than static images. The Speech Commands dataset is designed for voice recognition tasks, featuring audio recordings of spoken commands. Focusing on ten classes — digits from zero to nine — we converted the audio files into flattened one-dimensional arrays of Mel-frequency cepstral coefficients (MFCCs), which are common features used in speech recognition [44, 45]. Each MFCC feature matrix comprises 12 coefficients that capture the primary audio characteristics. For binary processing compatibility, these matrices were binarized so that non-positive elements were set to ‘0’ and positive elements to ‘1’. For detailed information on the preprocessing steps and the composition of the MFCC feature matrix, refer to the ‘Methods’ section and Supplementary Section 5.
Expanding our evaluation to the Speech Commands dataset, our results are detailed in Fig. 5, showcasing how the network adapts to audio recognition tasks. As anticipated, the classification accuracy for spoken digits increases with the number of neurons, peaking at about 68%, with results presented in Fig. 5(a). Similar to our findings with image-based datasets, the performance of our neuromorphic network excels beyond that of traditional linear classifiers both in the binarized (by about 19%) and non-binarized (by about 7%) MFCC feature arrays.
In examining the performance of our proposed neuromorphic network on the Speech Commands dataset, it’s beneficial to compare it against a well-established benchmark in speech recognition: the Hidden Markov Model with Gaussian Mixture Model (HMM-GMM) [39, 40]. This approach utilizes Hidden Markov Models to sequence audio data, where each state is characterized by a mixture of Gaussian distributions. Traditionally used for modeling audio sequences in speech recognition tasks, HMM-GMM operates on input data without undergoing binarization, providing a robust baseline with a reported accuracy of 40.9% across the full Speech Commands dataset [46] and 65.8% for the ten-class subset. Impressively, our binary neuromorphic network surpasses this result, achieving even greater accuracy on binarized data.
A notable observation from our study pertains to the application of the input signal densing technique. As depicted in Fig. 5(b), unlike with the MNIST dataset, adjusting the densing degree $s$ for the Speech Commands does not yield improvements in classification accuracy, regardless of the neuron count. The ineffectiveness of densing in this context can be attributed to the intrinsic data structure of Speech Commands, where the average filling of the initial binarized matrix is approximately 0.51, closely balancing the ‘0’ and ‘1’ signals. This balance is critical because the linear classifier used for post-processing in the network treats imbalances towards more 0’s or more 1’s with similar effectiveness.
For datasets characterized by a denser filling, where the majority of the input vector elements are active, achieving a balanced input suitable for optimal classifier performance can often be managed at the binarization stage by adjusting the threshold. Here, the threshold is the value at which pixels in the initial data vector are converted to binary values: pixels with intensities above this threshold are set to ‘1’, and those below or equal to the threshold are set to ‘0’. In the case of MNIST, where the average natural filling is significantly lower (around 0.19), achieving such balance using threshold adjustments alone is not feasible, thus highlighting the utility of the input signal densing technique.
It is somewhat presumptuous to specify an exact optimal filling degree for the input vector because in fact the parity should be achieved in the output signal entering the linear classifier, following its traversal through the hidden layer. Nevertheless, the balance of 0’s and 1’s in the input signal can serve as a rough guideline for approximating this parity.
Further details on the applicability of both threshold management and the input signal densing technique, particularly with the MNIST datasedt, are discussed in Supplementary Section S6. Supplementary Section S4 provides accuracy estimates for the Fashion-MNIST dataset, where a zero threshold in binarization (assigning a ‘1’ to any non-zero pixel) naturally leads to a filling that approaches parity. These insights underscore the boundaries and effectiveness of the input signal densing technique across different datasets.
@@ Discussion
To address the challenge of scaling polariton neural networks beyond current limitations, we propose several strategies. These strategies, detailed in Sec. S4 of the Supplementary Information, include tiling multiple spatial light modulators (SLMs), utilizing optical waveguides, implementing advanced micro-optics, and employing dynamic reconfiguration techniques. By using multiple SLMs placed adjacent to each other, one can create a larger composite grid that significantly increases the addressable area for the pump beams. This modular approach allows for the expansion of the network by simply adding more SLMs, with careful alignment and synchronization ensuring seamless operation across the entire grid. Optical waveguides can be integrated into the sample to direct light precisely to designated spots, overcoming the spatial limitations of free-space optics. These waveguides can be fabricated using advanced lithography techniques to create efficient light paths with minimal loss, enabling the creation of larger networks without increasing the physical footprint.
Incorporating advanced micro-optics, such as microlens arrays, allows for precise focusing and directing of pump beams. This enhances the density of pump spots within the available area, enabling more neurons to be addressed simultaneously. Techniques like diffractive optical elements can further optimize the beam distribution and intensity profile. Dynamic reconfiguration techniques enable a single SLM to sequentially address different regions of the sample at high speeds or to reproduce different configurations of the neural network on the same region. This time-multiplexing approach effectively increases the number of controllable neurons without increasing the physical size of the SLM or the sample, by dynamically reconfiguring the pump spots or the neural network fragments. Each of these approaches is designed to overcome spatial constraints and enhance the efficiency of pump beam delivery, thereby enabling the creation of larger and more complex polariton neural networks. For more detailed descriptions and technical aspects of these strategies, refer to the Supplementary Information.
The main advantage of the polaritonic part of the neural network is its unprecedented speed, with processing times of only a few tens of picoseconds required for the generation and evolution of polariton condensates in response to laser pump pulses. Delays caused by electronic components are inevitable for neural networks of any nature during data loading and result reading stages. However, all other stages in a polaritonic network can be realized without electronic components.
In each specific experiment, the input signal transformation masks, once set initially, remain unchanged throughout the experiment. This means that while the input signals change, the optical paths of individual signals remain constant. Among the operations mentioned, randomization and expansion are linear operations that can be described as matrix-vector multiplications. These operations can be optically implemented, as mentioned in [4, 5] and detailed in [47, 48]. An input light field representing the vector of input signals passes through an optical element that applies a spatially varying phase or amplitude modulation, effectively performing a matrix-vector multiplication at the speed of light.
The signal densing operation is more complex, but solutions exist for its optical implementation as well. Our studies on the dependence of neuron performance on signal pulse intensity show that, with other parameters fixed, the intensity of signals can vary within several folds while ensuring correct neuron switching and not compromising the isolation of dyads. In this mode, signal densing can be achieved by simply combining multiple pulses at a single point on the sample, resulting in a proportional increase in the intensity of the resulting pulse. Such a combination can be realized using optical components like beam splitters and combiners.
Another approach involves temporal multiplexing, where the steps of the signal densing process are spread out over time. This technique is similar to the dynamic reconfiguration strategy discussed for network scaling, where different steps of the process occur sequentially in time but within the same spatial region. By carefully timing these steps, we can ensure that the combined effect of the pulses is achieved without significant delay. Even with an increase in the resulting operation time to hundreds of picoseconds, this approach still maintains a substantial speed advantage over electronic neural networks.
@@
Transitioning to the aspects of practical realization of the proposed architecture, it becomes essential to consider the possible experimental realization of polariton condensate lattices designed above. Spatially-distributed systems comprising chains and lattices of interconnected quantum entities have gained recognition as platforms for information storage, transmission and processing, as well as simulators of complex phenomena [20, 49, 50, 51, 22]. Polariton lattices, a recent breakthrough in spatially-distributed quantum systems, are notable for their exceptional spatial coherence [30, 24, 29, 25]. This coherence significantly exceeds that of individual condensates and it facilitates phase locking of nodes across the entire lattice. This widespread phase synchronization could potentially enhance the nonlinearity of interactions essential for neuromorphic network operation, see, e. g., [4]. However, this scenario leads to the loss of a key advantage emphasized in our work: the feasibility of neuron response binarization would be significantly compromised, impacting computational speed and resource efficiency. To address this issue while retaining the advantage of polariton lattices, which is the optical controllability of connections between condensates, we propose ensuring pairwise interactions of condensates within the lattice by selectively severing non-contributing links. For this purpose, a variety of approaches exists.
@@
Our investigation into binary polariton neural networks has demonstrated their fundamental operational principles and potential for energy-efficient computing, suitable for certain applications where high speed and low power consumption are critical. Energy consumption and efficiency analysis of the proposed system are discussed in Sec. S5 of the Supplementary Information. While binary networks inherently trade off precision for efficiency, our results with a basic logistic regression model on the MNIST dataset have shown promising accuracy levels. Moreover, the proposed network architecture, while simple, suggests several avenues for enhancement that could address tasks requiring higher precision.
We propose that extending the network’s complexity through additional hidden layers could amplify its computational power. To address the challenge of integrating multiple hidden layers in a polaritonic neural network, the output signals from one hidden layer can be utilized as a template for generating input signals for the next layer. This approach may involve electronic components for signal regeneration or amplification, which, while enhancing accuracy, compromises the system’s computations speed advantage.
Alternatively, integrated optical waveguides for polaritons can link neurons across hidden layers. These waveguides utilize the bistability effect, where a polariton condensate in a low-intensity state is switched to a high-intensity state by an input trigger [8]. The hidden layer photoluminescence signals can serve as these triggers, effectively generating signals in neurons at the opposite ends of the waveguides in subsequent layers. This approach maintains the speed advantage by avoiding electronic mediation and capitalizes on the intrinsic properties of polaritons for rapid signal transmission. This approach, although requiring adjustments to the hidden layer geometry, offers a feasible and efficient solution for multi-layer integration in polaritonic neural networks.
Adjusting the network’s geometric layout from a square to a hexagonal lattice could further improve its performance by increasing the number of neurons each input signal can potentially activate, thus enhancing the structural nonlinearity of the network. Moreover, replacing the OR gate response with other types of logical operations, such as XOR gates, could add another layer of nonlinearity. An XOR gate could toggle the state of each neuron in response to paired inputs without changing the overall parity, introducing a dynamic component to the neuron’s response.
We also propose the idea of accommodating continuous input signals alongside binary outputs. This approach would utilize the varying intensities of signals, such as those from grayscale images, to modulate input signals, enabling a more nuanced response based on pixel brightness. By doing so, we could significantly enrich the input feature space without compromising the inherent advantages of binary systems, such as low memory usage and high processing speeds.
Finally, exploring controlled interactions between neurons — what is often considered a parasitic effect of crosstalk — could be harnessed beneficialthily. By fine-tuning the isolation and interaction among neuron pairs, we could potentially enhance the network’s robustness and precision.
@@
The network’s effectiveness was demonstrated using the MNIST dataset for handwritten digit recognition. Our network has not only shown competitive performance against existing systems, but also surpassed them when taking advantage of the original signal densing technique. The developed approach allowed the network to achieve a remarkable 97.5% classification accuracy, theoretically. Further validation was conducted on the Speech Commands dataset, which contains diverse and complex one-second audio clips of spoken words. This additional testing phase highlighted the adaptability and robustness of our architecture in processing intricate audio data and handling a variety of speech recognition tasks.
@@ Methods
$U_{\text{d}}(t,\mathbf{r})=\left(\frac{g_{1}+iR_{1}}{2(\gamma_{\text{R}}+R_{1}|\Psi(t,\mathbf{r})|^{2})}+G_{1}\right)P_{1}(t,\mathbf{r})$ (3)
@@
characterizes the complex effective potential arising from the excitation of polariton condensates in a dyad through non-resonant pulsed optical pumping with intensity $P_{1}(t,\mathbf{r})$. $g_{1}$ is the constant of interaction of polaritons with reservoir excitons, $R_{1}$ is the stimulated scattering rate from the reservoir to the condensate, $\gamma_{\text{R}}$ is the exciton decay rate. The parameter $G_{1}$ characterizes repulsion from the dark exciton reservoir, which also inevitably emerges within the pump spot.
@@
The third term in Eq. (2) is responsible for the potential, arising from the signal pulse of intensity $P_{2}(t,\mathbf{r})$:
$U_{\text{in}}(t,\mathbf{r})=\left(\frac{g_{2}+iR_{2}}{2(\gamma_{\text{R}}+R_{2}|\Psi(t,\mathbf{r})|^{2})}+G_{2}\right)P_{2}(t,\mathbf{r}).$ (4)
@@
In Ref. [30], it is suggested to use cross-circularly polarized pump $P_{2}(t,\mathbf{r})$ in relation to $P_{1}(t,\mathbf{r})$. This approach aims to reduce gain of the condensates in the dyad due to their overlapping with the signal. The parameters in Eq. (4) for $U_{\text{in}}(t,\mathbf{r})$ have the same meaning as those in Eq. (3) for $U_{\text{d}}(t,\mathbf{r})$. We take the pumps in a Gaussian form as
$P_{1}(t,\mathbf{r})=P_{10}\exp\left(-t^{2}/w_{\tau 1}^{2}\right)\sum_{j,k}\exp\left[-(x+jd)^{2}/2w_{1}^{2}+(y+kd)^{2}/2w_{1}^{2}\right],$ (5a)
$P_{2}(t,\mathbf{r})=\exp\left(-t^{2}/w_{\tau 2}^{2}\right)\sum_{j,k}P_{20}^{(j,k)}\exp\left\{-[x+(j+0.5)d]^{2}/2w_{2}^{2}+[y+(k+0.5)d]^{2}/2w_{2}^{2}\right\},$ (5b)
where $w_{\tau 1,\tau 2}$ and $w_{1,2}$ are durations and spatial widths of the pump pulses, $d$ is the distance between the centers of the pump spots. The amplitude $P_{10}$ is taken equal for all pulses across the lattice unless otherwise indicated. The amplitudes ${P_{20}^{(j,k)}=P_{20}}$ for ‘1’ elements of the signal lattice and ${P_{20}^{(j,k)}=0}$ for ‘0’ elements.
@@
The last term in Eq. (2), which differs our model from one in Ref. [30], characterizes the stationary barrier. The barrier $V(\mathbf{r})=F(\mathbf{r},V_{0},u,a)$, schematically shown in Fig. 2(c), is a function of a spatial coordinate, with the following parameters: $V_{0}$ is the height of the potential, $u$ is the width of the walls, and $a$ is the width of the gap in the wall, responsible for entrance of the input signal. Remarkably, the magnitude $V_{0}$ can be both real and imaginary, depending on the nature of the barrier.
Settings for simulation parameters
@@
For simulation, we take the following parameters suggested in Ref. [30]. The effective polariton mass is ${m^{*}=0.49\,\text{meV ps}^{2}\mu\text{m}^{-2}}$, the decay rates are ${\gamma=1/6\,\text{ps}^{-1}}$ and ${\gamma_{\text{R}}=0.05\,\text{ps}^{-1}}$, the interaction constants are ${\alpha=2.4\,\mu\text{eV}\,\mu\text{m}^{2}}$, ${g_{1}=0.8\alpha}$ and ${g_{2}=1.8\alpha}$, repulsion constants from dark reservoirs are $G_{1,2}=4g_{1,2}/\gamma_{\text{R}}$, the durations and spatial widths of the pump pulses are ${w_{\tau 1}=5}$ ps, ${w_{\tau 2}=8}$ ps and ${w_{1,2}=2.2\,\mu\text{m}}$, the distance between the pump spots is ${d=12\,\mu\text{m}}$. We also take the scattering rates as ${R_{1}=7\alpha}$ and ${R_{2}=0.7R_{1}}$.
Preprocessing of Speech Command audio data
The Speech Commands dataset [37, 38] comprises WAV audio files that are preprocessed through a defined sequence to make them suitable for the neuromorphic network analysis:
  • Audio Length Adjustment: While each audio file in the Speech Commands dataset is intended to be approximately one second long, some are shorter due to truncated recordings, often cutting off the very words they are meant to capture. We remove these incomplete files from our dataset to maintain consistency and quality, only using full-length recordings for our analysis.
  • MFCC Extraction: We compute MFCCs from each audio file, extracting 12 coefficients per frame across 24 frames. This configuration balances detailed audio analysis with computational efficiency, ensuring we capture essential speech characteristics effectively for each one-second clip. See Supplementary Section 5 for more details about MFCC feature matrix.
  • Binarization: The MFCC feature arrays are binarized. Non-positive values are set to 0 and positive values are set to 1.
  • Dataset Splitting: The processed data is split into training and testing sets using a 90:10 ratio, ensuring both effective training and thorough performance evaluation.
Acknowledgments
The support of Saint-Petersburg State University (research grant No. 122040800257-5), the state assignment in the field of scientific activity of the Ministry of Science and Higher Education of the Russian Federation (theme FZUN-2024-0019, state assignment of VlSU) and the Innovation Program for Quantum Science and Technology 2023ZD0300300 are acknowledged.
Authors’ contribution
E.S. devised the concept, performed simulations and analyzed the simulation results. A.K. initiated the study and guided the work. Both authors wrote the manuscript, discussed and contributed to the paper.
Conflict of interest
The authors declare that they have no conflict of interest.
Data availability
All data generated and analyzed during the current study are available from the corresponding author on reasonable request.